GBT Status Update

19 March 2001


The contractor and the NRAO suffered a setback last week in efforts to stabilize the movement in the azimuth track.  Last Monday, March 12, a segment of the wear strip was re-doweled with longer and larger diameter dowels (4 x 1.5 inch vs. 3 x 1 inch), compared to the previous experiment some weeks ago in which the dowels broke.  The dowels did not break this time, and initial measurements indicated that the track motion had been significantly reduced to about 0.020 inches in each direction as the wheels rolled back and forth over the segment.  However, a repeat measurement on Thursday showed that the motion had increased  to about 0.025 inches in each direction, suggesting that the holes were being elongated from use of the telescope.  Furthermore, about 20 hardened steel hold-down bolts at other locations around the track were found to be fractured. 

A significant reanalysis of the problem is underway now by the contractor, with the active participation of the NRAO.  The first scientific observations with the GBT on 24-26 March,  with a second installment on 3-4 April, will proceed as scheduled.  After those observations are finished, azimuth motion of the telescope will be suspended until the track problem can be further analyzed.  We will not be able to move the telescope for normal, astronomical commissioning observations during this period.  This will slow commissioning, and will probably delay the scheduling of further science projects until the problem is in hand.  Fortunately, we have had 1.5 months of extremely productive astronomical commissioning results, and have a lengthy list of system projects that can be pursued in the interregnum.  We will use the time as efficiently as possible to minimize the long-term impact of this delay.

Last week's commissioning observations again produced good results.  A few more pointing observations were obtained at S-band (~2 GHz) and a pointing model was fit to the all-sky data.  The total rms (vector sum of both axes) was about 8.5 arcsec.  These data had been taken on 9-10 March, then again on 12 March, and contained both nighttime and daytime data with sun on the structure. Control software for ephemeris tracking of solar system objects was installed and is ready for an on-object test this week.   Tests of IF stability across the fiber optic transmission line were conducted again with elevation motion only.  Over a large range of elevation, gain was stable to about 1 part in 10,000.  The VLBI backend was also tested last week.  One problem in the implementation of phase calibration was found and will be addressed.  Insofar as the backends are concerned, a VLBI run could be conducted now with the GBT.  Considerable software work remains to implement VLBA control of the GBT.

Metrology tests with the ground-based rangefinders continued last week.  For the first time, all 12 ground rangefinders were used at once, with each acquiring their spherical  retroreflector targets on the structure of the  GBT.  A complete data set was recorded, and will be used to confirm the structural model of the GBT.

P. R. Jewell

M. M. McKinnon